hydrogen gas is generated when acids come into contact with certain metals. When excess hydrocloric acid reacts with 2.5 g of Zn (the metal product is Zn^(2)), what volume of hydrogen gas is collected at a pressure of 0.93 atm and a temperature of 22\deg C?
1
Expert's answer
2019-09-06T04:16:59-0400

Zn+2HCl = ZnCl2+ H2

PV=nRT -> V=nRT/P

nH2=nZn=2.5/65.38=0.038 mole, P=0.98*101.3 *10^3= 99.274*10^3 Pa, T=295K

V= 0.038*8.31*295/(99.274*10^3)= 0.938*10^-3 m^3=0.938 l